1.Water flows at 750 m3/s down a rectangular channel made of finished concrete. The channel has a width of 4 m and a downward slope of 0.0002, and the water depth is 1.5 m at the control section A. Determine the distance x from A to where the depth becomes 1.25 m. Use delta y = 0.05 m and plot the surface profile. Take n = 0.012.


2. The rectangular channel is fitted with a 60 degrees or triangular weir plate. If the upstream depth of the water is 2 m. and the bottom of the weir plate is 1.6 m from the bottom of the channel. determine the volumetric flow of water prang over the weir. Take Cd = 0.72. 



1
Expert's answer
2022-01-20T10:08:26-0500

1.

"Q=\frac 1n(\frac{l(h-x)}{\Delta y})^{\frac 23}\sqrt I,"

"x=11~m."

2.

"Q=C\sqrt{hl^2\sin^2\alpha}=150~\frac{m^3}s."